OpenStudy (mynamachef):

Manipulating Equations E = mc^2 c = ? Possible Answers: E/m m/E sqrt(E/m) sqrt(m/E) E-m sqrt(E-m)

OpenStudy (osprey):

3rd one down MASS ENERGY EQUIVALENCE is the important thing about Einstein's equation http://perendis.webs.com

OpenStudy (jose4l):

E=m(c)^2 from this equation c = sqrt(E/m)

OpenStudy (osprey):

There is also the point that if you have something, possibly a photon, with the right amount of energy two particles can be generated "from the energy". This is part of what PARTICLE physics and some very very powerful accelerators are about (eg CERN, Europe, etc)